# Recent Advances in Predictive Modeling with Electronic Health Records

### v1

@@ Abstract
The development of electronic health records (EHR) systems has enabled the collection of a vast amount of digitized patient data. However, utilizing EHR data for predictive modeling presents several challenges due to its unique characteristics. With the advancements in machine learning techniques, deep learning has demonstrated its superiority in various applications, including healthcare. This survey systematically reviews recent advances in deep learning-based predictive models using EHR data. Specifically, we begin by introducing the background of EHR data and providing a mathematical definition of the predictive modeling task. We then categorize and summarize predictive deep models from multiple perspectives. Furthermore, we present benchmarks and toolkits relevant to predictive modeling in healthcare. Finally, we conclude this survey by discussing open challenges and suggesting promising directions for future research.
@@ 1 Introduction
Advancements in healthcare technology have sparked a revolution in patient information storage. An increasing number of health institutions and providers are now adopting electronic health record (EHR) systems to digitize patient information. This enables healthcare professionals in extracting valuable insights from EHR data, thereby supporting various aspects of healthcare,ranging from enhancing clinical decision-making to predicting diseases, enabling personalized medicine, driving quality improvement initiatives, and contributing to public health surveillance and epidemiological research.
@@
  • Clinical Explainability: Healthcare professionals often require predictive models to be explainable. Many advanced machine learning algorithms, while effective, may lack explainability, making it challenging for clinicians to understand, trust, and utilize the model results.
@@ 3 Existing Progress
RetainEX Kwon et al. (2018) TVCG Heart Failure Cataract HIRA
@@
TContextGGAN Xu et al. (2022) TKDE Heart Failure Sepsis Risk MIMIC
GRAM Choi et al. (2017) KDD Dignosis Prediction MIMIC, Private
KAME Ma et al. (2018b) CIKM Dignosis Prediction MIMIC, Private
PRIME Ma et al. (2018a) KDD Heart Failure, COPD kidney Disease Private
MedPath Ye et al. (2021a) WWW Heart Failure, COPD Kidney Disease Private
KerPrint Yang et al. (2023a) AAAI Disease Prediction MIMIC, Private
@@
DCMN Feng et al. (2019) ICDM Mortality, Cost MIMIC, Private
RAIM Xu et al. (2018) KDD Decompensation Length of Stay MIMIC
MUFASA Xu et al. (2021) AAAI Diagnoses MIMIC
AutoMed Cui et al. (2022) BIBM COPD, Amnesia Kidney Disease Private
AutoFM Cui et al. (2024) SDM Acute Renal Failure Diagnoses MIMIC
NNR Chen and Cai (2015) JMBE Diabetes Private
K-Boosted C5.0 Zhang et al. (2019) JHE Breast Cancer Private
@@
TWIN Das et al. (2023) KDD Breast Cancer, Lung Cancer Project Data Sphere
@@
LSAN Ye et al. (2020) CIKM Heart Failure COPD, Kidney Disease Private
ConCAD Huang and Ma (2021) ECMLPKDD Sleep Apnea Apnea-ECG, MIT-BIH PSG
@@
Table 1: Summarization of recent work on predictive modeling in healthcare. F1: modeling temporality; F2: modeling time irregularity; F3: incorporating extra knowledge; F4: interpretability; F5: modeling multimodal EHR data; F6: using automated machine learning (AutoML) techniques; and F7: addressing class imbalance issues.
@@ 3.1 Basic Deep Learning-based Predictive Models
Recurrent Neural Networks-based Models
@@
Transformer-based Models
@@ 3.3 Knowledge-Enhanced Predictive Modeling
A critical avenue for enhancing predictive models in the EHR domain lies in integrating additional external knowledge. Below we categorize the methods for leveraging external knowledge based on the type of knowledge being utilized.
Structured Knowledge
The GRAM framework Choi et al. (2017) utilizes a graph attention network to integrate hierarchical medical ontologies into the learning process of code representations. Building upon GRAM, KAME Ma et al. (2018b) advances this concept by embedding ontology information throughout the entire prediction process, thereby enriching the model’s contextual understanding. Beyond merely serving as an additional feature, external knowledge can also function as a regularization component. For instance, PRIME Ma et al. (2018a) generates a posterior regularization term from external knowledge, aligning the model’s predictions with established medical insights. This strategy simplifies incorporating external knowledge, making it broadly applicable across various EHR risk prediction frameworks. Besides graph attention, ConCAD Huang and Ma (2021) develops a cross attention mechanism to combine the deep representations with expert knowledge, which automatically fuses the knowledge by emphasizing the important parts based on each other synergistically.
@@
3.4 Preditive Modeling with Multimodal Data
@@
RAIM Xu et al. (2018) integrates waveform and vital signs as inputs into a multi-channel attention mechanism module. The module subsequently aggregates discrete clinical features into the generated embeddings for subsequent prediction tasks. Another approach, proposed by Yang et al. Yang and Wu (2021), focuses on fusing temporal clinical features, time-invariant data, and clinical notes through summation in the output stage. They explore various combinations of modality-specific encoders to achieve competitive performance.
A more recent work Wang et al. (2023) leverages an attention mechanism to obtain a weighted summation of multiple clinical modalities, including temporal clinical features, diagnosis, medication, clinical notes, and demographics. This approach aims to generate a meaningful multimodal Electronic Health Record (EHR) representation for both pretraining and task-specific fine-tuning.
Federated Learning
@@ 3.5 AutoML-based Predictive Modeling
MUFASA Xu et al. (2021) is the first work that applied the neural architecture search (NAS) technique to medical multimodal fusion. It jointly optimizes multimodal fusion strategies and modality-specific architectures. MUFASA delineates two types of blocks: modality-specific block, which only allows inputs from the same modality, and fusion block, which accepts inputs from both fusion architecture hidden states and modality-specific states. AutoMed Cui et al. (2022) employs a general Directed Acyclic Graph (DAG) framework, opting for a cell-based search where each cell consists of a sequentially ordered set of computation nodes. It uses the same search space design for modality encoding cells and fusion cells. AutoFM Cui et al. (2024), also focuses on modality-specific search and multimodal fusion search. It designs separate search spaces for static and sequential modalities. It incorporates modality interaction operations within these modality-specific search spaces to facilitate early interactions. A specialized set of simple fusion operations is devised to enable effective feature fusion.
3.6 Preditive Modeling with Imbalanced Classes
@@
Oversampling and Undersampling Techniques
@@
MIMIC-III Real 38,597 patients
Demographics, vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data
https://physionet.org/content/mimiciii/1.4/
MIMIC-IV Real 40,000+ patients
Demographics, vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data
https://physionet.org/content/mimiciv/2.2/
MIMIC-CXR Real 377,110 images 227,835 reports
Electronic health record data, images (chest radiographs), and natural language (free-text reports)
https://physionet.org/content/mimic-cxr/2.0.0/
eICU Real 200,000+ admissions
Vital sign measurements, care plan documentation, severity of illness measures, diagnosis information, treatment information, and more
https://physionet.org/content/mimic-cxr/2.0.0/
PPMI Real 2,230 patients
Subject characteristics, biospecimen, images, medical history, etc.
https://www.ppmi-info.org/
ADNI Real 2,775 patients
Subject characteristics, genetic data, images, medical history, neuropathology, etc.
https://adni.loni.usc.edu/
Apnea-ECG Real 70 recordings
Subject characteristics, electrocardiogram
https://physionet.org/content/apnea-ecg/1.0.0/
MIT-BIH PSG Real 18 recordings
Subject characteristics, electrocardiogram, electroencephalography, electrooculography, electromyography, etc.
https://physionet.org/content/slpdb/1.0.0/
SHHS Real 6,441 patients
Subject characteristics, electrocardiogram, electroencephalography, electrooculography, electromyography, airflow, etc.
https://sleepdata.org/datasets/shhs
Newcastle-Accel Real 28 patients
Subject characteristics, acceleration, polysomnography.
https://zenodo.org/records/1160410##.YLqiSC1h1qt
Sleep-Accel Real 31 patients
Acceleration, heart rate, steps.
https://physionet.org/content/sleep-accel/1.0.0/
EMRBOTS Synthetic 100,000 patients
Patients’ admissions, demographics, socioeconomics, labs, medications, etc.
http://www.emrbots.org/
Project Data Sphere Real 242 studies
Data provider, sponsor, study phase, linked data, tumor type, access, etc.
https://www.projectdatasphere.org
@@
3.7 Interpretable Preditive Modeling
Interpretability is an important property of health risk prediction models because it is related to the life and death of patients when they are applied in real life. Interpretability helps make sure that the doctors and patients can understand the reasoning of the predictive model and trust the risk prediction or treatment recommendation of the predictive model. Existing interpretability mechanisms designed in the health risk prediction models can be mainly divided into the following three categories.
Attention-based Interpretation
The first type of interpretability method uses attention weights to explain the importance of different diagnosis codes or visits. This method has been widely used for RNN-based methods Choi et al. (2016b); Ma et al. (2017); Bai et al. (2018); Ma et al. (2018c) as we have mentioned above. Attention mechanisms can also be incorporated into Transformer-based predictive models such as HiTANet Luo et al. (2020) and LSAN Ye et al. (2020). Take LSAN as an example for illustration. It introduces a hierarchical attention mechanism that assigns flexible attention weights to different diagnosis codes by their relevance to corresponding diseases in the diagnosis code level and pays greater attention to visits with higher relevance in the visit level. Such design provides a fine-grained interpretability. The attention weights in the hierarchy of diagnosis codes and visits can be interpreted into which symptoms and visits are paid more attention to for health risk prediction.
Personalized Knowledge Graph-based Interpretation
@@ 3.7 Interpretable Preditive Modeling
Medical Text-based Explicit Interpretation
@@
Uncertainty-based Interpretation
@@ 4 Benchmarks and Toolkits
Publicly Available Healthcare Databases
@@
Private Healthcare Databases
@@ 5.2 Data Scarcity/Sparsity
5.3 Pre-training Across Multiple Data Sources

### v2

@@ Abstract
The development of electronic health records (EHR) systems has enabled the collection of a vast amount of digitized patient data. However, utilizing EHR data for predictive modeling presents several challenges due to its unique characteristics. With the advancements in machine learning techniques, deep learning has demonstrated its superiority in various applications, including healthcare. This survey systematically reviews recent advances in deep learning-based predictive models using EHR data. Specifically, we introduce the background of EHR data and provide a mathematical definition of the predictive modeling task. We then categorize and summarize predictive deep models from multiple perspectives. Furthermore, we present benchmarks and toolkits relevant to predictive modeling in healthcare. Finally, we conclude this survey by discussing open challenges and suggesting promising directions for future research.
@@ 1 Introduction
Advancements in healthcare technology have sparked a revolution in patient information storage. Many health institutions and providers are now adopting electronic health record (EHR) systems to digitize patient information. This enables healthcare professionals to extract valuable insights from EHR data, thereby supporting various aspects of healthcare, ranging from enhancing clinical decision-making to predicting diseases, enabling personalized medicine, driving quality improvement initiatives, and contributing to public health surveillance and epidemiological research.
@@
  • Clinical Explainability: Healthcare professionals often require predictive models to be explainable. While effective, many advanced machine learning algorithms may lack explainability, making it challenging for clinicians to understand, trust, and utilize the model results.
@@ 3 Existing Progress
GRAM Choi et al. (2017) KDD Dignosis Prediction MIMIC, Private
@@
RetainEX Kwon et al. (2018) TVCG Heart Failure Cataract HIRA
KAME Ma et al. (2018b) CIKM Dignosis Prediction MIMIC, Private
PRIME Ma et al. (2018a) KDD Heart Failure, COPD kidney Disease Private
RAIM Xu et al. (2018) KDD Decompensation Length of Stay MIMIC
DCMN Feng et al. (2019) ICDM Mortality, Cost MIMIC, Private
K-Boosted C5.0 Zhang et al. (2019) JHE Breast Cancer Private
@@
LSAN Ye et al. (2020) CIKM Heart Failure COPD, Kidney Disease Private
MedPath Ye et al. (2021a) WWW Heart Failure, COPD Kidney Disease Private
MUFASA Xu et al. (2021) AAAI Diagnoses MIMIC
@@
ConCAD Huang and Ma (2021) ECMLPKDD Sleep Apnea Apnea-ECG, MIT-BIH PSG
TContextGGAN Xu et al. (2022) TKDE Heart Failure Sepsis Risk MIMIC
AutoMed Cui et al. (2022) BIBM COPD, Amnesia Kidney Disease Private
@@
KerPrint Yang et al. (2023a) AAAI Disease Prediction MIMIC, Private
TWIN Das et al. (2023) KDD Breast Cancer, Lung Cancer Project Data Sphere
AutoFM Cui et al. (2024) SDM Acute Renal Failure Diagnoses MIMIC
Table 1: Summarization of recent work on predictive modeling in healthcare (in publication years). F1: modeling temporality; F2: modeling time irregularity; F3: incorporating extra knowledge; F4: interpretability; F5: modeling multimodal EHR data; F6: using automated machine learning (AutoML) techniques; and F7: addressing class imbalance issues.
@@ 3.1 Basic Deep Learning-Based Predictive Models
Recurrent Neural Networks-Based Models.
@@
Transformer-Based Models.
@@ 3.3 Knowledge-Enhanced Predictive Modeling
A critical avenue for enhancing predictive models in the EHR domain lies in integrating additional external knowledge. Below, we categorize the methods for leveraging external knowledge based on the type of knowledge being utilized.
Structured Knowledge.
GRAM Choi et al. (2017) utilizes a graph attention network to integrate hierarchical medical ontologies into the learning process of code representations. Building upon GRAM, KAME Ma et al. (2018b) advances this concept by embedding ontology information throughout the entire prediction process, thereby enriching the model’s contextual understanding. Beyond merely serving as an additional feature, external knowledge can also function as a regularization component. For instance, PRIME Ma et al. (2018a) generates a posterior regularization term from external knowledge, aligning the model’s predictions with established medical insights. This strategy simplifies incorporating external knowledge, making it broadly applicable across various EHR risk prediction frameworks. Besides graph attention, ConCAD Huang and Ma (2021) develops a cross-attention mechanism to combine deep representations with expert knowledge.
@@
3.4 Predictive Modeling with Multimodal Data
@@
RAIM Xu et al. (2018) integrates waveform and vital signs as inputs into a multi-channel attention mechanism module. The module subsequently aggregates discrete clinical features into the generated embeddings for subsequent prediction tasks. Another approach, proposed by Yang et al. Yang and Wu (2021), focuses on fusing temporal clinical features, time-invariant data, and clinical notes through summation in the output stage. The researchers explore various combinations of modality-specific encoders to achieve competitive prediction performance.
A more recent contribution by Wang et al. Wang et al. (2023) leverages an attention mechanism to obtain a weighted summation of multiple clinical modalities, including temporal clinical features, diagnosis, medication, clinical notes, and demographics. This approach aims to generate a meaningful multimodal Electronic Health Record (EHR) representation for both pretraining and task-specific fine-tuning.
Federated Learning.
@@ 3.5 AutoML-Based Predictive Modeling
MUFASA Xu et al. (2021) is the first work that applied the neural architecture search (NAS) technique to medical multimodal fusion. It jointly optimizes multimodal fusion strategies and modality-specific architectures. MUFASA delineates two types of blocks: modality-specific block, which only allows inputs from the same modality, and fusion block, which accepts inputs from both fusion architecture hidden states and modality-specific states. AutoMed Cui et al. (2022) employs a general Directed Acyclic Graph (DAG) framework, opting for a cell-based search where each cell consists of a sequentially ordered set of computation nodes. It uses the same search space design for modality encoding cells and fusion cells. Another noteworthy work, AutoFM Cui et al. (2024), also focuses on modality-specific search and multimodal fusion search. This approach designs separate search spaces for static and sequential modalities. Furthermore, it incorporates modality interaction operations within these modality-specific search spaces to facilitate early interactions. During the fusion search phase, a specialized set of simple fusion operations is devised to enable effective feature fusion.
3.6 Predictive Modeling with Imbalanced Classes
@@
Oversampling and Undersampling Techniques.
@@
MIMIC-III Real 38,597 patients Demographics, vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data https://physionet.org/content/mimiciii/1.4/
MIMIC-IV Real 40,000+ patients Demographics, vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data https://physionet.org/content/mimiciv/2.2/
MIMIC-CXR Real 377,110 images 227,835 reports Electronic health record data, images (chest radiographs), and natural language (free-text reports) https://physionet.org/content/mimic-cxr/2.0.0/
eICU Real 200,000+ admissions Vital sign measurements, care plan documentation, severity of illness measures, diagnosis information, treatment information, and more https://physionet.org/content/mimic-cxr/2.0.0/
PPMI Real 2,230 patients Subject characteristics, biospecimen, images, medical history, etc. https://www.ppmi-info.org/
ADNI Real 2,775 patients Subject characteristics, genetic data, images, medical history, neuropathology, etc. https://adni.loni.usc.edu/
Apnea-ECG Real 70 recordings Subject characteristics, electrocardiogram https://physionet.org/content/apnea-ecg/1.0.0/
MIT-BIH PSG Real 18 recordings Subject characteristics, electrocardiogram, electroencephalography, electrooculography, electromyography, etc. https://physionet.org/content/slpdb/1.0.0/
SHHS Real 6,441 patients Subject characteristics, electrocardiogram, electroencephalography, electrooculography, electromyography, airflow, etc. https://sleepdata.org/datasets/shhs
Newcastle-Accel Real 28 patients Subject characteristics, acceleration, polysomnography. https://zenodo.org/records/1160410##.YLqiSC1h1qt
Sleep-Accel Real 31 patients Acceleration, heart rate, steps. https://physionet.org/content/sleep-accel/1.0.0/
EMRBOTS Synthetic 100,000 patients Patients’ admissions, demographics, socioeconomics, labs, medications, etc. http://www.emrbots.org/
Project Data Sphere Real 242 studies Data provider, sponsor, study phase, linked data, tumor type, access, etc. https://www.projectdatasphere.org
@@
3.7 Interpretable Predictive Modeling
Interpretability is an important property of health risk prediction models because it is related to the life and death of patients when they are applied in real life. Interpretability helps doctors and patients understand the reasoning behind the predictive model and trust its risk prediction or treatment recommendation. Existing interpretability mechanisms designed in the health risk prediction models can be divided into three categories.
Attention-Based Interpretation.
The first type of interpretability method utilizes attention weights to explain the importance of different diagnosis codes or visits. This method has been widely used for RNN-based methods Choi et al. (2016b); Ma et al. (2017); Bai et al. (2018); Ma et al. (2018c) as we have mentioned above. Attention mechanisms can also be incorporated into Transformer-based predictive models such as HiTANet Luo et al. (2020) and LSAN Ye et al. (2020). Take LSAN as an example for illustration. It introduces a hierarchical attention mechanism that assigns flexible attention weights to different diagnosis codes by their relevance to corresponding diseases in the diagnosis code level and pays greater attention to visits with higher relevance in the visit level. Such design provides a fine-grained interpretability. The attention weights in the hierarchy of diagnosis codes and visits can be interpreted into which symptoms and visits are paid more attention to for health risk prediction.
Personalized Knowledge Graph-Based Interpretation.
@@ 3.7 Interpretable Predictive Modeling
Medical Text-Based Explicit Interpretation.
@@
Uncertainty-Based Interpretation.
@@ 4 Benchmarks and Toolkits
Publicly Available Healthcare Databases.
@@
Private Healthcare Databases.
@@ 5.2 Data Scarcity/Sparsity
5.3 Pretraining across Multiple Data Sources
@@ 5.4 Federated Training for Foundation Models
Acknowledgements
This work is partially supported by the National Science Foundation under Grant No. 2238275 and the National Institutes of Health under Grant No. R01AG077016.